# MSTN: Fast and Efficient Multivariate Time Series Model

### v1

@@ 4.2.1 Long Term Forecasting
The proposed MSTN model was evaluated on eight widely-used benchmark datasets: Weather, Traffic, Electricity, ILI, and four variants of ETT (ETTh1, ETTh2, ETTm1, ETTm2) [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64]. These datasets are standard in the field and have been extensively used in prior studies, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime] and iTransformer [liu2024itransformerinvertedtransformerseffective]. Notably, large-scale datasets such as Weather, Traffic, and Electricity contain a significantly higher number of time series, making their results more stable and less prone to overfitting compared to smaller datasets. We evaluate all methods across multiple prediction horizons $H\in{96,192,336,720}$ for standard benchmarks and $H\in{24,36,48,60}$ for the ILI dataset.
4.2.2 Imputation
The performance of imputation is assessed on benchmark datasets from the ETT [zhou2021informer], Electricity [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather] domains, which commonly encounter missing data in real-world applications. Following the protocol established by TimesNet [wu2023timesnettemporal2dvariationmodeling], we evaluated performance under random masking ratios of $12.5\%,25\%,37.5\%,50\%$}. The evaluation includes recent SOTA baselines such as GPT2(3) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
@@ 4.2.3 Classification
For classification tasks, we utilized ten multivariate benchmark datasets from the UEA time series classification archive [bagnall2018ueamultivariatetimeseries]. These datasets encompass diverse applications including gesture and action recognition, heartbeat-based medical diagnosis, and other real-world scenarios. The evaluation compares with recent SOTA baselines such as GPT2(6) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], DLinear [zeng2022transformerseffectivetimeseries], ETSFormer [woo2022etsformer], FEDformer [zhou2022fedformer], and Informer [zhou2021informer].
4.2.4 SOTA Baselines for Generalizability Study
To rigorously evaluate the generalizability of the proposed MSTN framework, we conducted cross-domain benchmarking across seven publicly available international datasets. The evaluation encompasses datasets from diverse domains: human safety, with the Boubézoul et al. [BOUBEZOUL2020105577] (France) fall event dataset and the Rodegast et al. [DARUS33012023] (Germany) simulator for risky riding behaviour and collisions; general human activity, using UCI-HAR [har240] (Italy) and PAMAP2 [pamap2_physical_activity_monitoring_231] (USA) to assess robustness in physical activity recognition; animal behavior monitoring via the ActBe-Calf dataset [DISSANAYAKE2025111462] (Ireland) for welfare assessment; and mechanical systems, including MetroPT-3 [metropt3] (Portugal) for predictive maintenance and the NASA Turbofan Engine [NasaTurboFan] (USA) for sensor degradation forecasting. This selection tests model transferability across healthcare, activity recognition, agricultural technology, and industrial monitoring. All models were evaluated under a consistent five-fold cross-validation framework to ensure a robust and fair comparison, including recent architectures such as TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
4.3 Training and Model Configuration
@@
The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1+cu118 and trained on an NVIDIA T400 GPU (4 GB VRAM). The architecture processes variable-length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128→64 filters). Training uses AdamW [kingma2017adammethodstochasticoptimization] optimizer with learning rate $3\times 10^{-4}$, batch size 64, and task-specific objectives (Focal Loss for classification, MSE for forecasting and imputation). Models train for up to 100 epochs with early stopping based on validation performance for task. Dataset-specific configurations include 96-step lookback for Traffic data and 50-step sequences for physical therapy datasets, with min-max feature normalization applied throughout.
@@ 7 Author Contributions
Sumit S Shevtekar led the research design, developed the MSTN architecture, implemented all software, conducted experiments across classification, forecasting and imputation tasks, performed data analysis and visualization, and wrote and edited the original draft. Chandresh K Maurya contributed to methodology design, experimental validation, and manuscript writing, review and editing. Gourab Sil participated in manuscript review and editing. All authors reviewed and approved the final manuscript.
8 Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 4.3.2 Long Term Forecasting
The proposed MSTN model is evaluated for the long-term time series forecasting task. We evaluated all methods across multiple prediction horizons $H\in\{96,192,336,720\}$ for standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI dataset. The forecast performance is quantified using MSE and MAE. The evaluation includes the following SOTA baselines: EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
@@ 4.3.3 Classification
The performance of classification tasks is evaluated using the standard metric of classification accuracy. The evaluation compares with recent SOTA baselines such as GPT2(6) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], DLinear [zeng2022transformerseffectivetimeseries], ETSFormer [woo2022etsformer], FEDformer [zhou2022fedformer], and Informer [zhou2021informer].
4.3.4 Generalizability Study
To evaluate generalizability between domains of the proposed MSTN framework, we benchmarked against baselines (e.g. TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64]) on seven publicly available international datasets spanning healthcare, activity recognition, agricultural technology, and industrial monitoring. All evaluations utilized a consistent five-fold cross-validation framework to ensure a robust and fair comparison.
4.4 Experimental Setup
4.4.1 Training and Model Configuration
@@
The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1 and trained on an NVIDIA T400 GPU (4 GB VRAM). The architecture processes variable length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128 $\to$ 64 filters). Training uses AdamW [kingma2017adammethodstochasticoptimization] optimizer with learning rate $3\times 10^{-4}$, batch size 64, and task-specific objectives (Focal Loss for classification, MSE for forecasting and imputation). Models train for up to 100 epochs with early stopping based on validation performance. Data preprocessing employs min-max normalization to standardize feature distributions. For forecasting and imputation benchmarks, we utilize a fixed input lookback window of $L=96$ (except $L=36$ for ILI) across all prediction horizons, whereas for classification tasks involving variable length sequences (e.g., JapaneseVowels), inputs are standardized to the maximum sequence length (e.g., $L=29$) to facilitate consistent batch processing.
4.4.2 Evaluation Metrics
The performance of forecasting and imputation tasks is evaluated using two primary regression metrics that measure the discrepancy between predicted values and ground truth. For forecasting tasks, these metrics quantify the accuracy of future predictions, while for imputation tasks they assess the quality of missing value reconstruction. The performance is evaluated using:
Mean Squared Error: $\quad\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(Y_{i}-\hat{Y}_{i})^{2}$ (15)
Mean Absolute Error: $\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}|Y_{i}-\hat{Y}_{i}|$ (16)
where $Y,\hat{Y}\in\mathbb{R}^{H\times C}$ are the ground truth and prediction results of the future with $H$ time points and $C$ channels. $Y_{i}$ denotes the $i$-th future time point. The performance of classification tasks is evaluated using four standard classification metrics that measure different aspects of predictive performance, particularly important for handling potential class imbalance in the datasets. The classification performance is evaluated using:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (17)
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (18)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (19)
$\text{F1-Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
where TP denotes True Positives, TN True Negatives, FP False Positives, and FN False Negatives.
@@ 7 Author Contributions
Sumit S Shevtekar led the research design, developed the architecture, implemented all software, conducted experiments, performed data analysis and visualization, and wrote and edited the original draft. Chandresh K Maurya contributed to experimental validation, manuscript writing, review, and editing